OpenStudy (anonymous):

Find the smaller solution for 3/5x^2−7=21

OpenStudy (anonymous):

x^2 = 3/140

OpenStudy (anonymous):

what ?

OpenStudy (anonymous):

what is wrong?

OpenStudy (anonymous):

what are you saying !?

OpenStudy (anonymous):

x to the power of 2= 3 divided by 140

OpenStudy (anonymous):

thats wrong :(

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

yes :/

OpenStudy (anonymous):

there are other simplifications... that is just one of them

OpenStudy (anonymous):

thanks either way -.-

OpenStudy (asnaseer):

is this your question?\[\frac{3}{5}x^2-7=21\]

OpenStudy (anonymous):

yes !

OpenStudy (asnaseer):

ok, then we get:\[\frac{3}{5}x^2=21+7=28\]\[3x^2=5*28=140\]\[x^2=\frac{140}{3}\]

OpenStudy (anonymous):

but you have to find the square root

OpenStudy (asnaseer):

so then you just take the square root to get x = plus/minus some value

OpenStudy (asnaseer):

also, what do you mean by smaller solution? e.g. if answer is \(\pm10\), then do you mean -10 as this is less than +10?

OpenStudy (anonymous):

yea whatever the answer is i just put negative and it should be right

OpenStudy (asnaseer):

ok, so you should be able to find the answer now from what I have given you above

OpenStudy (anonymous):

can you just tell me please am being timed

OpenStudy (zepp):

Okay, from there \[x^2 = 140/3\] You know that \[a^2 = b\] could also be \[a =\sqrt {b}\] Now do the same thing with it and it will give \[x = \pm\sqrt {140/3}\] The smallest value would be \[x = - \sqrt {140/3}\] the only thing you have to do is to simplify it